### v1

# OpenHuEval: Evaluating Large Language Model on Hungarian Specifics

Abstract
We introduce OpenHuEval, the first benchmark for LLMs focusing on the Hungarian language and specifics. OpenHuEval is constructed from a vast collection of Hungarian-specific materials sourced from multiple origins. In the construction, we incorporated the latest design principles for evaluating LLMs, such as using real user queries from the internet, emphasizing the assessment of LLMs’ generative capabilities, and employing LLM-as-judge to enhance the multidimensionality and accuracy of evaluations. Ultimately, OpenHuEval encompasses eight Hungarian-specific dimensions, featuring five tasks and 3953 questions. Consequently, OpenHuEval provides the comprehensive, in-depth, and scientifically accurate assessment of LLM performance in the context of the Hungarian language and its specifics. We evaluated current mainstream LLMs, including both traditional LLMs and recently developed Large Reasoning Models. The results demonstrate the significant necessity for evaluation and model optimization tailored to the Hungarian language and specifics. We also established the framework for analyzing the thinking processes of LRMs with OpenHuEval, revealing intrinsic patterns and mechanisms of these models in non-English languages, with Hungarian serving as a representative example. We will release OpenHuEval at https://github.com/opendatalab/OpenHuEval .
OpenHuEval: Evaluating Large Language Model on Hungarian Specifics
Haote Yang1thanks: Equal contribution., Xingjian Wei [^1] footnotemark: 1, Jiang Wu [^1] footnotemark: 1thanks: Project lead., Noémi Ligeti-Nagy211footnotemark: 1, Jiaxing Sun311footnotemark: 1, Yinfan Wang411footnotemark: 1, Zijian Győző Yang211footnotemark: 1, Junyuan Gao511footnotemark: 1, Jingchao Wang6, Bowen Jiang7, Shasha Wang1, Nanjun Yu6, Zihao Zhang8, Shixin Hong9, Hongwei Liu1, Wei Li1, Songyang Zhang1, Dahua Lin1,10, Lijun Wu1, Gábor Prószéky2, Conghui He1thanks: Corresponding author. 1Shanghai Artificial Intelligence Laboratory, 2HUN-REN Hungarian Research Centre for Linguistics, 3Wuhan University, 4Shanghai University, 5University of Chinese Academy of Sciences, 6East China Normal University, 7Peking University, 8Harbin Institute of Technology, 9Tsinghua University, 10Chinese University of Hong Kong Correspondence: heconghui@pjlab.org.cn
1 Introduction
Recent advancements in Large Language Models (LLMs) Jaech et al. (2024); Team (2024a)represent significant strides toward artificial general intelligence (AGI). However, notable performance gaps remain between English and other languages in both language-agnostic tasks (e.g., math reasoning, code generation) Huang et al. (2025); Zhang et al. (2024) and language-specific tasks (e.g., idiom usage, cultural understanding) Naous et al. (2023); Sun et al. (2024); Liu et al. (2024c), posing challenges to global AI deployment and equitable development. The disparities in cross-lingual performance arise mainly from two factors: First, the training data, particularly the pre-training data, is heavily skewed toward English. Second, while English evaluation benchmarks are advanced and rapidly evolving, non-English benchmarks are underdeveloped, particularly for language-specific features, limiting the identification of shortcomings in non-English languages and leading to their neglect in research.
This paper focuses on the evaluation of Hungarian language and specifics. Hungarian is spoken by around 14 million people worldwide. Research on the Hungarian language not only improves the user experience for Hungarian speakers but also offers valuable insights for similar studies in other languages and regions. Existing Hungarian evaluation datasets are largely translations of English ones, missing essential Hungary-specific elements such as language nuances, culture, history, and regional context, which are key for Hungarian users. Among the existing evaluation datasets, HuLU Ligeti-Nagy et al. (2024) is the key benchmark for Hungarian language understanding, but its focus on multiple-choice and true/false questions limits its ability to evaluate broader LLM capabilities, such as language generation, open-domain Q&A, reasoning and instruction-following.
Benchmark
Real user
query
Self-awareness
evaluation
Proverb
Reasoning
Generative task
& llm-as-judge
Hungarian
Lang
Comprehensive
Hu-specific
WildBenchLin et al. (2024)
SimpleQAWei et al. (2024),
ChineseSimpleQAHe et al. (2024)
MAPSLiu et al. (2024c)
MARC, MMMLU et al in Lai et al. (2023)
BenchMAXHuang et al. (2025)
MILQANovák et al. (2023)
HuLULigeti-Nagy et al. (2024)
OpenHuEval (ours)
Table 1: Comparison of related benchmarks.
Figure 1: Overview of OpenHuEval.
To address this gap, we introduces OpenHuEval, the first evaluation benchmark for LLMs focused on Hungarian language and specifics. The comparison of OpenHuEval with the existing related benchmarks is shown in Table 1. Overall, OpenHuEval has two main distinguishing features:
1) Hungarian-Specific Inspired by Liu et al. (2024b); Sun et al. (2024), we propose eight distinct Hungarian-specific dimensions (see §2.1), covering a variety of scenarios that users may encounter when querying in Hungarian. Guided by these dimensions, we collected a vast amount of Hungarian specific material from multiple sources and used this to construct the corresponding evaluation tasks.
2) Keeping up with the latest advances in LLM evaluation Significant progress has been made in LLM evaluation, with query sources shifting from manual or rule-based constructions to real-world internet questions Lin et al. (2024), enhancing practical relevance. Question formats evolved from multiple-choice to open-ended Q&A Wei et al. (2024), better reflecting actual usage. Evaluation methods transitioned from rule-based approaches to LLM-as-judge and subjective assessments, improving accuracy and objectivity Li et al. (2024). However, these advancements primarily apply to English datasets and not Hungarian. Thus, when creating OpenHuEval, we incorporated these principles and methodologies from English evaluations.
Based on OpenHuEval, we evaluated the performance of mainstream LLMs on Hungarian language and specifics. We compared the performance differences of these models on the typical datasets of OpenHuEval with corresponding datasets in other languages. The results indicate a significant necessity for evaluation and model optimization specifically for Hungarian language and specifics.
Moreover, Large Reasoning Models (LRMs), like o1, mark a new direction in LLM development. Through extensive reasoning, self-reflective negation, and exploring multiple reasoning paths, they greatly improve reasoning abilities, adhering to the test-time scale law—a crucial step toward AGI. Recent studies Wang et al. (2025) have analyzed these reasoning processes, offering insights for optimization, but have largely focused on English-language contexts, neglecting Hungarian language and specific scenarios. Building on OpenHuEval, we developed the framework for dissecting the reasoning processes of LRMs. Using Hungarian as the example, we uncovered intrinsic patterns of the representative LRMs in non-English languages. These findings provide valuable insights for the research community to further advance the development of LRMs.
In summary, the contributions of this paper include the following three points:
- We developed OpenHuEval, the first benchmark for LLMs focusing on the Hungarian language and specifics. OpenHuEval incorporates the latest design principles for evaluating LLMs, such as using real user queries from the internet, emphasizing the assessment of LLMs’ generative capabilities, and employing LLM-as-judge to enhance the multidimensionality and accuracy of evaluations.
- We conducted the comprehensive evaluation of current mainstream LLMs, including traditional LLMs and recently developed LRMs. The results highlight the significant necessity for evaluation and model optimization tailored to Hungarian language and specifics.
- We established the framework for analyzing the thinking processes of the cutting-edge LRMs, revealing the intrinsic patterns and mechanisms of these models in non-English languages, with Hungarian as a representative.
Figure 2: Examples of OpenHuEval. The original text is in black, while the translation into English is in blue. In HuMatchingFIB and HuStandardFIB, the blank is highlighted.
HuSpecificDim
Definition #Question
$\mathcal{L}$) Basic knowledge of the Hungarian language and Hungarian proverbs and sayings 1333
$\mathcal{H}$) Historical events and historical development of Hungary 324
Life, Culture, and Custom($\mathcal{LCC}$)
Religion, rituals, culture, holidays, and the daily life of Hungarians 622
Education and Profession($\mathcal{EP}$)
Education system in Hungary and related professions 282
Geography and Place($\mathcal{GP}$)
Geographical knowledge of Hungary, cities, and locations 165
$\mathcal{F}$) Famous figures of Hungary 409
Politics, Policy and Law($\mathcal{PPL}$)
Politics, policies, and laws of Hungary 485
Business and $\mathcal{BF}$)
Business and finance in Hungary 333
Table 2: The Hungarian-Specific Dimensions (HuSpecificDim).
2 OpenHuEval
OpenHuEval is the benchmark specifically designed to evaluate the performance of LLM in handling Hungarian language and specifics. The overview of OpenHuEval is in Figure 1. Examples are shown in Figure 2.
2.1 Hungarian-specific dimensions and OpenHuEval tasks
Inspired by Liu et al. (2024b); Sun et al. (2024), OpenHuEval encompasses eight Hungarian-Specific Dimensions (HuSpecificDim), as shown in Table 2: Language ($\mathcal{L}$), History ($\mathcal{H}$), Life, Culture, and Customs ($\mathcal{LCC}$), Education and Profession ($\mathcal{EP}$), Geography and Place ($\mathcal{GP}$), Figure ($\mathcal{F}$), Politics, Policy, and Law ($\mathcal{PPL}$), and Business and Finance ($\mathcal{BF}$). These dimensions comprehensively cover a wide range of scenarios encountered by users when utilizing Hungarian as the query language. As a result, they enable the systematic evaluation of the performance of LLMs in tasks related to the Hungarian.
Task
HuSpecificDim
Judge
Question type
#Question
HuWildBench $\mathcal{LCC}$, $\mathcal{EP}$, $\mathcal{PPL}$, $\mathcal{BF}$ llm,checklist OE 1154
HuSimpleQA $\mathcal{L}$$\mathcal{H}$,$\mathcal{LCC}$,$\mathcal{EP}$,$\mathcal{GP}$$\mathcal{F}$,$\mathcal{PPL}$$\mathcal{BF}$ llm OE 1293
HuProverbRea $\mathcal{L}$ rule,llm 2CQ/OE 1135
HuMatchingFIB $\mathcal{L}$, $\mathcal{H}$ rule Matching Filling-in-Blank 278
HuStandardFIB $\mathcal{L}$, $\mathcal{H}$ rule,similarity matching Standard Filling-in-Blank 93
Table 3: Tasks of OpenHuEval
Bearing the above HuSpecificDim in mind, the first step in building OpenHuEval involves gathering corpora rich in Hungarian specifics. Inspired by Liu et al. (2024c); Li et al. (2023); He et al. (2024); Arora et al. (2024), we collected data from sources such as Hungarian proverbs, exam questions, forums, and Wikipedia. Through processes including filtering, refinement, construction, and quality assurance, we developed a total of five evaluation tasks comprising 3953 questions in total, as detailed in Table 3. The subsequent sections of this chapter will introduce these tasks and their corresponding datasets in detail.
2.2 Hungarian WildBench
Task Introduction: The Hungarian WildBench (HuWildBench) task aims to evaluate the performance of LLMs in answering various questions arising from the everyday lives of Hungarians. All questions are sourced from Hungary’s well-known forum website [^1] https://www.gyakorikerdesek.hu/, which is similar to https://www.quora.com/ for English-speaking world. (hereinafter referred to as “g13k” for brevity). These questions cover a wide range of topics, including cultural customs, education, tourism, legal regulations, and business and finance, thus reflect real-life issues encountered by Hungarians. Examples of HuWildBench questions are shown in Figure 2 and Table 15. The queries in HuWildBench are user-generated content, therefore their linguistic expressions and question formats tend to be less formal than the structured and polished written language. This poses the realistic challenge for LLMs, as they must adapt to such informal and spontaneous language style. The construction of HuWildBench is detailed in Appendix B.
Metric and judge: We use the WB-Score Lin et al. (2024) as evaluation metric for HuWildBench, which is calculated in the following manner: We have simultaneously developped the customized checklist for each question during the process of creating each question, to assist the LLM judge in evaluating the responses. Examples of these checklists can be seen in Table 15. Following Lin et al. (2024), GPT-4o is used as the judge model, which evaluates the quality of each response based on the checklist and provides detailed strengths and weaknesses before assigning a score from 1 to 10. The definition of scores is shown in Table 6 and the judge prompt is detailed in Appendix B.5. Different from Lin et al. (2024), our final scores are calculated as the average of all test sample scores, with each score multiplied by 10.
2.3 Hungarian SimpleQA
Task Introduction: Hungarian SimpleQA (HuSimpleQA) is designed to assess the ability of LLMs to answer short, fact-seeking questions related to Hungary. Inspired by Wei et al. (2024) and He et al. (2024), we constructed HuSimpleQA based on Hungarian Wikipedia [^2] https://hu.wikipedia.org/, with the following key characteristics. Hungarian: The questions in HuSimpleQA are in Hungarian, and they focus on facts specifically related to Hungary. Diverse: The questions in HuSimpleQA cover the eight Hungary-specific dimensions proposed in §2.1. High-quality: The construction process of HuSimpleQA (in Appendix C) includes comprehensive and strict quality control procedures, ensuring the quality and accuracy of the questions. Static: Similar to SimpleQA, the answers to the questions in HuSimpleQA do not change over time, ensuring that the dataset remains evergreen. Easy-to-evaluate: The questions and answers in HuSimpleQA are short and concise, making them ideal for evaluation through LLMs. The examples of HuSimpleQA are shown in Figure 2 and Table 16. The construction of HuSimpleQA is detailed in Appendix C.
Metric and Judge: Following Wei et al. (2024), we use GPT-4o as a judge to categorize the responses of the LLM to HuSimpleQA into three classes: CORRECT, INCORRECT or NOT_ATTEMPTED. The judge prompt can be found in Appendix C.7, Figure 27.
2.4 Hungarian Proverb Reasoning
Task Introduction: Hungarian Proverb Reasoning (HuProverbRea), which consists of the collection of Hungarian proverbs, idioms, abbreviations, is a task that requires the LLM to understand and reason the meaning of Hungarian proverbs in a specific context. As shown by the examples in Figure 2, LLM is provided with a context in which a Hungarian proverb is used, accompanied by a question: “What does the speaker mean by the saying?”. Then, the LLM is tasked with discerning the speaker’s true intention, either by selecting the correct interpretation from two provided options (2CQ setting), or by directly articulating the speaker’s intended meaning (OE setting). The construction of HuProverbRea is detailed in Appendix D.
Metric and judge: For the 2CQ setting, we simply measure the correct ratio of candidate LLMs. For the OE setting, we adopt GPT-4o as judge to decide if the answer is acceptable. We provide the original proverb, its context and the English explanation of the proverb as references when judging OE responses. Detailed prompt templates are listed in Appendix D.
2.5 Hungarian Matching and Standard Filling-in-Blank
Task Introduction: Hungarian Matching Fill-in-the-Blank (HuMatchingFIB) is a task where key terms in a text are removed, and a pool of candidate words or phrases is provided. This pool includes both correct answers and distractors. The task requires the LLM to choose the most suitable words from the pool to fill in the blanks, thus restoring the full meaning of the text. The example is shown in Figure  2 and Figure  33. HuMatchingFIB effectively tests the LLM’s abilities in understanding information, reasoning within context, and differentiating correct answers from distractors.
In contrast, Hungarian Standard Fill-in-the-Blank (HuStandardFIB) also involves filling in blanks but does not offer a candidate pool. Instead, the model must rely on its internal knowledge and the provided context to complete the text. The examples are shown in Figure 34. Consequently, HuStandardFIB evaluates the LLM’s overall ability to recall knowledge and reason within context.
The constructions of HuMatchingFIB and HuStandardFIB are detailed in Appendix E.
Metric and Judge: In our inference prompts, we explicitly instruct the LLM to generate responses in a specified format. Additionally, we have established a set of well-defined rules to evaluate the correctness of the LLM’s answers for each blank. The detailed format requirements and judgment criteria can be found in the Appendix E.
For both HuMatchingFIB and HuStandardFIB, we evaluate performance at two levels: blank-level and question-level accuracy. Specifically, Acc_b (Blank-level Accuracy) measures the proportion of blanks that the model answers correctly across all questions. On the other hand, Acc_q (Question-level Accuracy) evaluates the proportion of questions that the model answers entirely correctly. A question is only considered correct if all its associated blanks are answered accurately.
3 Experiments and Analysis
3.1 Experimental setup
We evaluated the currently mainstream LLMs, including GPT-4o Hurst et al. (2024), GPT-4o mini [^3] We used the gpt-4o-2024-11-20 version for GPT-4o and the gpt-4o-mini-2024-07-18 version for GPT-4o-mini., Deepseek-V3 Liu et al. (2024a), Qwen2.5-Instruct Yang et al. (2024), and Llama-3.1-Instruct Dubey et al. (2024), as well as the latest Large Reasoning Models (LRMs) such as OpenAI o1-mini Jaech et al. (2024), QwQ-32B-Preview Team (2024b) (abbreviated as QwQ in following text), and Deepseek-R1 Team (2024a). Detailed specifications of these models are provided in Table 4.
We used OpenCompass [^4] https://github.com/open-compass/opencompass in all our experiments. For traditional instruction-based LLMs, we adopted OpenCompass’s default settings for the maximum output length. For Large Reasoning Models, we set the output length to 8192 to ensure sufficient space for reasoning process and to produce a complete final answer, avoiding premature output truncation. For OpenAI models (GPT series and o1-mini), we used their official API with settings following OpenCompass’s default configuration. For Deepseek-V3 and Deepseek-R1, due to the high usage volume of Deepseek’s official API causing instability, we used equivalent API services provided by Alibaba Cloud [^5] https://cn.aliyun.com/ and Silicon Valley Flow [^6] https://siliconflow.cn/. The settings followed OpenCompass’s configurations, with the temperature set to 0.7. For other models in Table 4, we performed inference locally with NVIDIA A100 GPUs, using LMDeploy [^7] https://github.com/InternLM/lmdeploy as the inference backend. The settings followed OpenCompass’s default configuration (Temperature = 1e-6, top_k = 1).
  Model Size
Reasoning
Model
Open-
source
Inference Method
GPT-4o - N N Official API
GPT-4o-mini - N N Official API
Deepseek-V3 - N Y Alibaba Cloud and SiliconFlow API
Qwen2.5-Instruct 7B,72B N Y Local GPU
Llama-3.1-Instruct 8B,70B N Y Local GPU
o1-mini - Y N Official API
QwQ 32B Y Y Local GPU
Deepseek-R1 - Y Y
Alibaba Cloud and
SiliconFlow API
Table 4: LLMs evaluated in our experiments.
3.2 Overall performance
  HuWildBench HuSimpleQA HuProverbRea HuMatchingFIB HuStandardFIB
Model WBScore Acc Acc. (OE) Acc. (2CQ) B acc. Q acc. B acc. Q acc.
  GPT-4o 81.09 50.3 89.16 95.51 77.78 43.88 57.36 15.05
GPT-4o-mini 74.19 25.56 84.67 92.16 55.68 19.78 35.08 7.53
QwQ 58.02 9.09 67.49 84.23 38.65 12.23 6.05 0
Deepseek-R1 82.96 34.58 82.29 91.72 80.87 47.12 61.76 17.2
Deepseek-V3 78.42 32.71 83.26 92.51 68.87 39.93 51.44 9.68
Llama-3.1-Instruct-70B 61.78 35.99 80.18 93.83 59.56 24.46 40.99 6.45
Llama-3.1-Instruct-8B 53.62 15.2 63.35 73.48 5.74 0.72 16.64 1.08
o1-mini 76.43 15.8 77.44 87.67 60.83 17.63 45.25 13.98
Qwen2.5-Instruct-72B 74.05 14.9 77.8 90.22 63.8 24.1 32.32 8.6
Qwen2.5-Instruct-7B 42.01 5.22 50.48 67.05 31.88 1.08 7.43 0
Table 5: Overall performance of 10 LLMs on OpenHuEval. The first, second, and third place in each metric are marked with red, green, and blue text, respectively. In the FIB task evaluation metric, B represents the blank level, and Q represents the question level.
The overall performance of all LLMs on OpenHuEval is presented in Table 5. It can be observed that across a total of five tasks, Deepseek-R1 ranks first in three tasks and achieves top-tier performance in the other two tasks. GPT-4o ranks first in two tasks and second in the remaining three tasks. These results demonstrate the exceptional performance of the two models in Hungary-specific tasks.
Open-source models vs Closed-source models: Among open-source models, Deepseek-R1 stands out, while Deepseek-V3 also demonstrates strong overall performance, ranking highly across all tasks. Llama-3.1-Instruct-70B achieved impressive scores of 93.83% in the HuProverbRea-2CQ task and 36% in the HuSimpleQA task, ranking second only to the closed-source model GPT-4o. This highlights the growing potential of open-source models, led by Deepseek-R1, which are increasingly showing capabilities comparable to closed-source models in Hungarian language tasks.
Traditional LLMs vs. Large Reasoning Models: We compared Traditional LLMs and LRMs within the same series. Across five tasks, Deepseek-R1 consistently outperforms Deepseek-V3 in four of them. Specifically, in the HuMatchingFIB task, Deepseek-R1 achieves relative improvements of 12% at the blank level and 7.19% at the question level compared to Deepseek-V3. Similarly, for the HuStandardFIB task, it achieves gains of 10.32% (blank level) and 7.52% (question level). Although Deepseek-R1 performs slightly worse than Deepseek-V3 on the HuProverbRea task, the performance gap is less than 1%. Considering that both Deepseek-R1 and Deepseek-V3 are based on the same pretrained model, the significantly stronger performance of Deepseek-R1 on the OpenHuEval benchmark demonstrates the effectiveness of LRMs architectures in Hungarian language tasks and domain-specific scenarios. This result underscores the potential of LRMs as a key avenue of exploration in advancing Artificial General Intelligence (AGI).
Model size: From the results, models with larger parameter sizes perform better on OpenHuEval. For example, GPT-4o, Llama-3.1-Instruct-70B, and Qwen2.5-Instruct-72B outperform their smaller counterparts in the same series (GPT-4o-mini, Llama-3.1-Instruct-7B, and Qwen2.5-Instruct-7B) across all tasks.
Figure 3: Comparison of model performance on OpenHuEval and similar datasets, highlighting that most LLMs experience rank changes.
3.3 Comparison with Existing Benchmarks
We compare the LLM’s performance rankings on the datasets in OpenHuEval and the existing similar datasets:
HuWildBench vs WildBench Lin et al. (2024): As shown in Figure 3(a), 70% LLMs experienced ranking changes. Detailed results can be found in Table  7.
HuSimpleQA vs SimpleQA Wei et al. (2024): As shown in Figure 3(b), 70% LLMs experienced ranking changes. Detailed results can be found in Table 9.
HuProverbRea vs MAPS: HuProverbRea was constructed with reference to MAPS Liu et al. (2024c), which is the proverb reasoning dataset comprising six subsets, each corresponding to a different language: English (en), German (de), Russian (ru), Bengali (bn), Mandarin Chinese (zh), and Indonesian (id). We compared the model performance rankings on HuProverbRea and each subset of MAPS. As shown in Figure 3(c), the percentage of LLMs with ranking changes were: 70% for en, 80% for zh, 90% for ru & bn, and 100% for id & de. Detailed results can be found in Table 10 and Table 11.
These results underscore the importance of evaluating LLMs on Hungarian proverbs and Hungarian-specific questions, highlighting the need for targeted optimization of models to better handle language-specific proverbs and cultural nuances across diverse languages.
4 Framework for Analyzing the Thinking Process of LRM
When responding to the user’s query, the LRM’s response typically consists of two parts: the thinking process and the answer. [^8] In DeepSeek-V1, the thinking process and answer are enclosed within ¡think¿ ¡/think¿ and ¡answer¿ ¡/answer¿ tags, respectively. We developped the framework for the in-depth analysis of the LRM’s thinking process on OpenHuEval. For LRM, we chose Deepseek-R1 and QwQ, as these are the only two models with accessible reasoning processes.
4.1 Task Selection
Among the OpenHuEval tasks, we selected HuSimpleQA and HuMatchingFIB as our subjects of study. Unlike recent work Wang et al. (2025), which focuses solely on math reasoning datasets, the two tasks we selected each have distinctive characteristics: HuSimpleQA assesses the LLM’s ability to recall and retrieve Hungarian-specific knowledge, as well as its awareness of its own knowledge boundaries. HuMatchingFIB involves questions where multiple competitive blanks exist within the same problem, requiring the model to carefully choose which answers to fill in. Therefore, analyzing these tasks allows us to explore the reasoning mechanisms of LRMs in both broader and deeper contexts, providing the research community with more valuable conclusions and insights.
4.2 Analysis on HuSimpleQA
Figure 4: Method for analyzing the LRM’s thinking process on HuSimpleQA.
Method: Following Wang et al. (2025), the LRM’s thinking process can be broken down into “thoughts”. A “thought” refers to the intermediate cognitive step output by a LRM during its thinking process. Throughout the thinking process, the LRM transitions between multiple thoughts, which are typically separated by reflective phrases such as “Alternative”, “Várni”(wait). The examples of the thoughts and the transitions can be found in Figure 40 and Figure 41. Thoughts can be further classified as “correct” or “incorrect”: reasoning along correct thoughts leads to CORRECT responses, while incorrect thoughts result in INCORRECT.
The LRM’s responses to HuSimpleQA have been judged as CORRECT, INCORRECT, or NOT_ATTEMPTED. Then we used GPT-4o to split the thinking process into thoughts (see the prompt in Figure 37 and Figure 38). We evaluated the correctness of each thought (see the prompt in Figure 39), with examples provided Figure 40 and Figure 41 in Appendix F.
Figure 5: Average number of tokens and thoughts generated per response on HuSimpleQA.
Efficiency of thinking process: We measure the length of the process (in terms of token count) and the number of thoughts under the three evaluation outcomes of the responses, as shown in Figure 5. The results indicate that both the reasoning length (in tokens) and the thought count were generally shorter for Deepseek-R1 compared to QwQ. Considering that Deepseek-R1 performs better than QwQ on the HuSimpleQA task, it suggests that Deepseek-R1 achieves its superior performance with relatively lower reasoning overhead.
Confidence in thinking process: For Deepseek-R1, the reasoning length and thought count showed no significant differences across the three types of evaluation outcomes. In contrast, for QwQ, the length and the number of thoughts were significantly higher in the NOT_ATTEMPTED cases compared to the other two types. This observation suggests that, compared to Deepseek-R1, QwQ is less “confident”, which tends to repeatedly attempt generating answers when faced with uncertainty and is more inclined to abstain from answering altogether.
Figure 6: The distribution of thought correctness ratio in INCORRECT and NOT_ATTEPTED responses. Red lines for HuSimpleQA and blue lines for MATH500-Hard from Wang et al. (2025).
Correct thoughts in INCORRECT responses: In HuSimpleQA, we analyzed the ratio of INCORRECT and NOT_ATTEMPTED responses from Deepseek-R1 and QwQ that contained correct thoughts. We compared these results with those from MATH500-Hard as reported in Wang et al. (2025), as shown in Figure 6. In the math reasoning task (MATH500-Hard), the significant portion of LRM’s ultimately incorrect responses still included correct thoughts: 72% for Deepseek-R1 and 78% for QwQ contained at least one correct thought. However, in HuSimpleQA, the ratio of INCORRECT responses containing correct thoughts dropped to 46% for Deepseek-R1 and 42% for QwQ. These findings suggest that the reasoning processes of LRMs differ significantly between memory-intensive tasks (HuSimpleQA) and reasoning-intensive tasks (MATH500-Hard), highlighting the need for targeted analysis and optimization.
Figure 7: Method for analyzing the LRM’s thinking process on HuMatchingFIB.
Figure 8: Example for splitting DeepSeek-R1’s thinking process into segments and categorizing these segments on HuMatchingFIB.
4.3 Analysis on HuMatchingFIB
Method: Unlike HuSimpleQA where each query contains only one question, HuMatchingFIB involves multiple competitive blanks within the same question that need to be filled (see the example in Figure 2 and Figure 33). Based on extensive case studies, we have developed the analytical method to facilitate further in-depth analysis, as illustrated in Figure 7.
Initially, we split the thinking process of LRMs into several segments, categorizing each as introduction, reasoning, review, or summary. Typically, each thinking process begins with an introduction segment, includes several reasoning segments and some review segments in the middle, and concludes with a summary segment. In each reasoning segment, the LRM addresses one or more blanks through reasoning. During the review segment, there is usually a reflection on the completed blanks, with potential revisions to the answers for earlier blanks. Examples of these segments can be found in Figure 8, and detailed definitions are provided in Table 12.
We further tagged the reasoning segments according to the following four dimensions: (Dim1) correctness: Are the answers in this reasoning segment correct? (Dim2) complexity: In this reasoning segment, does the LRM simply assert the answer, or does it involve more complex reasoning? (Dim3) scope: Does this reasoning segment focuses on a single blank, modifies previous blanks, or addresses multiple blanks? (Dim4) language transfer: Does the LRM switch languages within this reasoning segment? The details of the tagging can be found in Appendix F.2.2. Examples of the tagging results can be found in Figure 47 and Figure 48.
Figure 9: Tagging results of LRM’s reasoning segments in the correctness and complexity dimensions.
Simple Assertion or Complex Thought? As shown in Table 13, the reasoning segments can be categorized into two classes. The first is referred to as Simple Assertion, where LRM directly provides the answer to the blank. The second type is termed Complex Thought, where the segment involves repeated thinking, logical reasoning, hypothesis validation, or other complex processes. Examples can be found in Figure 47 and 48.
We analyzed the correctness of the reasoning segments for both Simple Assertion and Complex Thought (tagging results of dim1), as shown in Figure 9. By comparing the statistical distributions of the Deepseek-R1 and QwQ models, we noted the following: Firstly, the proportion of Simple Assertion segments that are not Completely Correct is quite low (3% for Deepseek-R1 and 6% for QwQ), indicating that both models achieve high accuracy when resolving blanks through Simple Assertion. This suggests that the models’ calibration is relatively reliable, implying that the models “know what they know”. Secondly, the proportion of Simple Assertion segments that are Completely Correct is significantly higher for Deepseek-R1 at 49%, compared to only 31% for QwQ. This difference reflects the performance disparity between the two models in the thinking processes. Thirdly, the correctness for Complex Thought is notably lower than for Simple Assertion, and both models show a higher proportion of cases where no conclusion is reached for the Complex Thought tag. This indicates that when faced with unfamiliar problems, the LRM could actively extend the reasoning and analysis process.
Explicit Translation Insertion (ETI): We observed that in some reasoning segments, when faced with the problem in Hungarian, the LLM first translates the key phrase of the original question into English and then proceeds with analysis and reasoning based on this translation. For example, “… Erőteljes #3# és a költői #4# gazdag használata jellemzi. This translates to "It is characterized by strong #3# and rich use of poetic #4#." …”. We refer to this phenomenon as Explicit Translation Insertion (ETI). Statistical analysis shows that ETI occurs in 5.49% of DeepSeek-R1’s reasoning segments, while for QwQ, the proportion is 16.77%. This indicates that QwQ is relatively weaker in handling non-English inputs, tending to translate first and then reason, which is consistent with previous research findings on cross-lingual Chain-of-Thought (CoT) in LLMs Sun et al. (2024).
5 Conclusion
In this paper, we constructed the first benchmark for LLMs focusing on the Hungarian language and its specifics. The results highlight the significant need for evaluation and model optimization tailored to Hungarian language and specifics. We also developed the framework for analyzing the thinking processes of the cutting-edge LRMs. Our work not only advances LLM technology in Hungarian but also provides valuable insights for studying languages of other countries and regions.
6 Limitation
Given the rapid advancements in English evaluation datasets, this research represents only an initial step towards bridging the gap between Hungarian and English evaluation resources. Currently, evaluation datasets for less-resourced languages like Hungarian still lag behind their English counterparts in terms of depth and breadth. Moving forward, we plan to closely monitor developments in English evaluation methodologies, continually refining and enhancing evaluation techniques and datasets for low-resource languages to reduce this disparity.
Moreover, as the field of LLMs evolves rapidly, many promising models, especially those tailored for low-resource languages, remain under-evaluated. Our future goal is to establish a vibrant OpenHuEval community that will regularly update evaluation results for the latest models. This will ensure comprehensive and up-to-date assessments, fostering the optimization and development of models in the low-resource language domain.
7 Ethical Consideration
This work involved human annotation. For all annotators, we explicitly informed them about the use of the data and required them to ensure that the questions included in OpenHuEval do not involve any social bias, ethical issues or privacy concerns during the annotation process.
Acknowledgments
This research was supported by Shanghai Artificial Intelligence Laboratory.
Appendix A Related Works
A.1 Hungarian benchmarks for LLMs from Translation
Directly translating an existing English evaluation dataset into non-English is an effective and straightforward method for constructing non-English evaluation datasets. Many existing multilingual evaluation datasets are constructed using this approach, such as those described in Ahuja et al. (2023); Huang et al. (2024). Among these, quite a few evaluation datasets include Hungarian as one of the languages, such as Hu et al. (2020) and Adelani et al. (2023). Lai et al. (2023) utilized GPT-3.5 to translate datasets such as ARC Clark et al. (2018), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), and HellaSwag Zellers et al. (2019) into Hungarian. Belebele Bandarkar et al. (2023) includes Hungarian as part of its multilingual parallel corpus reading comprehension evaluation dataset. BenchMaX Huang et al. (2025) is a multilingual parallel corpus evaluation dataset focused on comprehensively assessing LLMs’ capabilities across various generative tasks.
These translation-based parallel corpus evaluation datasets provide a comprehensive assessment of LLMs’ language-agnostic capabilities in Hungarian, including areas like world knowledge, mathematical reasoning, logical reasoning, and code generation. However, they overlook the unique characteristics and linguistic features of Hungarian, such as language nuances, culture, history, and regional context, which are critical for Hungarian users.
A.2 Language specific benchmarks for LLMs
There has been extensive research focusing on evaluating the unique features and capabilities of languages Liu et al. (2024b). Some studies have constructed benchmarks similar to MMLU Hendrycks et al. (2020) by collecting exam questions specific to various countries, such as CMMLU Li et al. (2023), IndoMMLU Koto et al. (2023), ArabicMMLU Koto et al. (2024), KMMLU Son et al. (2024), and TurkishMMLU Yüksel et al. (2024). Other studies have built evaluation datasets by crawling material and user queries from internet forums, filtering for queries related to linguistic and cultural features, such as the benchmark in Naous et al. (2023) and CaLMQA Arora et al. (2024). Some works manually construct evaluation datasets that emphasize linguistic and cultural characteristics, such as BLEnD Myung et al. (2024), CHARM Sun et al. (2024), and HuLU Ligeti-Nagy et al. (2024). Others adopt a “LLM generation combined with human expert modification and review” method to construct culturally characteristic evaluation datasets, such as MAPS Liu et al. (2024c), mCSQA Sakai et al. (2024), and ChineseSimpleQA He et al. (2024). Although Hungarian is rarely covered in these works, they provide crucial inspiration and approaches for our work.
Among these works, MILQA Novák et al. (2023) is the Hungarian question answering benchmark created mainly following the SQuAD Rajpurkar et al. (2016). HuLU Ligeti-Nagy et al. (2024) is a comprehensive Hungarian evaluation benchmark kit  [^9] https://hulu.nytud.hu/ that includes a total of seven tasks and corresponding datasets. Of the seven tasks, four are constructed by translating existing English evaluation datasets, and three are manually created based on native Hungarian corpora. However, HuLU only supports multiple-choice and true/false questions, which limits its ability to assess broader LLM capabilities such as language generation, open-domain Q&A, reasoning, and instruction-following.
Appendix B HuWildBench
B.1 Overall Construction Pipeline of HuWildBench
Figure 10: Constuction of HuWildBench.
The construction of HuWildBench contains the following steps(see Figure 10):
(1) Crawling: All user queries on the g13k website [^10] https://www.gyakorikerdesek.hu/ are systematically categorized into a multi-level tag system, which consists of 27 primary tags and 231 secondary tags. We manually reviewed all the secondary tags and selected 37 of them that contain a higher number of questions related to Hungary, such as népszokások (folk customs), egészségügyi-ellátások (healthcare services), and rezsi (overheads). We then crawled user queries under these 37 secondary tags, with a query date range from January 1, 2019, to August 31, 2024, resulting in the dataset $Q_{crawl}$(approximately 523K queries).
(2) Filtering for Hungary-specific content: Although these 37 secondary tags are closely related to Hungary, many of the questions still do not focus on Hungary-specific topics. Therefore, we used GPT-4o to classify the questions in $Q_{crawl}$ (detailed prompt in Figure 13), resulting in a subset of approximately 260K questions, $Q_{specific}$
(3) Deduplication: To ensure the diversity of questions, we performed deduplication on the Hungary-related questions within each secondary tag. The detailed process is outlined in Appendix B.2. After deduplication, the number of user questions in $Q_{dedup}$ was approximately 134K.
(4) Automatic high-quality question filtering: To ensure that only high-quality questions are extracted from the question pool, we designed a comparative-based high-quality question filtering strategy, as detailed in Appendix B.3. After filtering, the resulting set $Q_{high\_quality}$ contained around 2K questions.
(5) Manual filtering: We hired a group of Hungarian native speakers to further manually review the questions in $Q_{high\_quality}$. Only questions that met the following two criteria were retained: First, the question should be Hungary-specific and closely related to Hungary. Second, the question must be harmless, meaning it does not contain inappropriate content such as pornography, violence, politics, or taboo topics specific to Hungary. The final set $Q_{manual\_check}$ consists of 1731 questions.
(6) Checklist construction: Based on WildBench Lin et al. (2024), we constructed a checklist for each question. The purpose of the checklist is to assist the LLM judger in evaluating the answers. Each item in the checklist queries a specific aspect of the answer to a question. An example of the checklist can be found in Table 15, and the detailed construction method is provided in Appendix B.4. To ensure the relevance of the checklist items to the questions, we hired a Hungarian native speaker to review the checklist for quality, filtering out non-compliant items and performing deduplication. The filtering criterion was whether the item was suitable as an evaluation dimension for the model’s response. To ensure the reliability of the LLM-as-judge, we filtered out user questions with fewer than 8 checklist items. The final set $Q_{checklist}$ contains 1154 questions. In the end, we obtained 1154 user questions along with their corresponding checklists.
B.2 Deduplication of similar questions
Since there are similar questions in the results obtained in the previous step, we design a method to remove similar ones. Specifically, we first use the SentenceTransformer Osváth et al. (2023) [^11] https://github.com/UKPLab/sentence-transformers model to extract the embedding vector of each question. Then, we calculate the cosine similarity between the embedding of each two questions, and choose a threshold between [0.15-0.25] according to the number of questions under each secondary tag.The larger the number of problems, the larger the threshold. Finally, one of the questions whose similarity is less than the threshold is removed, ensuring that the similarities between all questions are greater than the threshold.
B.3 Automatic high-quality question filtering
In order to automate the filtering of high-quality sample pots, we constructed the prompt that allows the GPT-4o to select the two best Hungarian questions out of the five based on the criteria of linguistic complexity, Hungarian relevance, common-sense accuracy, context-dependence, answer diversity, ambiguity, reasoning requirements, socio-ethical considerations, format diversity, and breadth of knowledge and outputs their indexes in JSON format to output their indexes, as shown in Figure 14. Specifically, we first set the criteria for high-quality questions in Prompt. Then we ask GPT-4o to compare the input questions based on the criteria. In order to mitigate the occurrence of some high-quality questions being eliminated prematurely (or vice versa) when all the questions in the same batch are of high quality, we follow the following 3 rules when filtering the high-quality questions: (1) filter 2 high-quality questions from 5 questions at a time, instead of filtering 1 high-quality question directly from 2 questions. (2) use the Swiss system mechanism instead of the knockout mechanism. In each screening round, each question can win in the current round as long as it ensures that it wins in two comparisons, and it will not be eliminated directly because of a failure in one comparison. (3) Our question screening strategy eliminates 65% of the questions in each round, in order to ensure that each secondary label has a sufficient number of high-quality questions. We conducted different elimination rounds for questions under different labels, and finally got about 2K questions. Finally, in order to validate the strategy of high-quality question screening, we manually checked about 200 5-option-2 results, and the pass rate was more than 80%, which proved the effectiveness of the present strategy. The final constructed HuWildBench is shown in Table 15.
B.4 Checklist construction
In the process of building the Checklist, we mainly use large language models to generate it. In order to ensure the diversity of the Checklist and make the judge model can better evaluate the quality of the answers, here we use two non-open source LLM GPT-4o and Claude-3.5, each model generates a list of length 3-5. then we merge the two Checklists into one final Checklist. Checklists are then merged into a final Checklist. ultimately, each problem has a length of 6-10 and a Checklist. The details of our designed Prompt are shown in Figure 15 and the final constructed partial Checklist is shown in Table 15.
B.5 LLM-as-judge
  Score Definition
Score 1-2 The response is very poor and does not make sense at all.
Score 3-4 The response is poor and does not help the user solve the problem meaningfully.
Score 5-6 The response is fair but has issues (e.g., factual errors, hallucinations, missing key information).
Score 7-8 The response is good but could be improved.
Score 9-10 The response is perfect and provides helpful information to solve the problem.
Table 6: Definition of WB-Scores.
Following WB-Score Lin et al. (2024), we use GPT-4o as the judge model to evaluate the score of responses generated by LLM. The prompt used in judge is shown in Figure 16.
B.6 Comparing the Performance Ranking on WildBench and HuWildBench
We compare the performance rankings of LLM on HuWildBench and WildBench Lin et al. (2024), as shown in Table  7.
  Rank WildBench HuWildBench
  1 Deepseek-V3 Deepseek-R1(↑3)
2 GPT-4o GPT-4o(-)
3 o1-mini Deepseek-V3(↓2)
4 Deepseek-R1 o1-mini(↓1)
5 Qwen2.5-Instruct-72B GPT-4o-mini(↑1)
6 GPT-4o-mini Qwen2.5-Instruct-72B(↓1)
7 Llama-3.1-Instruct-70B Llama-3.1-Instruct-70B(-)
8 QwQ QwQ(-)
9 Qwen2.5-Instruct-7B Llama-3.1-Instruct-8B(↑1)
10 Llama-3.1-Instruct-8B Qwen2.5-Instruct-7B(↓1)
Table 7: Performance rankings on WildBench and HuWildBench.
Appendix C HuSimpleQA
Figure 11: Constuction of HuSimpleQA.
C.1 Construction pipeline of HuSimpleQA
The construction of HuSimpleQA consists of the following steps, as shown in Figure 11:
(1) Obtaining corpora rich in Hungary-specific facts and knowledge: First, we chose the Hungarian Wikipedia website  [^12] https://hu.wikipedia.org/ as the source of corpus material for question construction. We crawled all the entry pages and extracted their content. Next, we used GPT-4o to classify whether the entries were Hungary-specific, with the prompts detailed in Appendix C.2, Figure 17.  [^13] We did not classify all the pages but instead randomly selected pages until we reached 8K Hungary-specific entries, at which point we stopped. We then used GPT to extract key information from the content of these entries suitable for question-answering. The extraction prompts are detailed in Appendix C.2, Figure 18. An example of the extracted key information is shown in Figure 19. As a result, we obtained Hungary-specific key information covering the eight distinct dimensions, totaling 4428 pieces of information.
(2) Generating questions and corresponding answers: We used the GPT-4o model to generate open-ended questions and corresponding answers based on the Hungary-specific key information obtained in the previous step. The prompt used is detailed in Appendix C.3, Figure 20. In this step, we generated a total of 9424 questions based on 4K entries. We then classified the generated questions according to the eight Hungary-specific dimensions outlined in Section 2.1, using GPT-4o with the prompt detailed in Appendix C.3, Figure 21.
(3) Automatic quality checking of questions: To ensure the quality of the questions, we used GPT-4o to check and filter the generated questions. We set the following four criteria, retaining only those questions that met all four standards (the corresponding prompt is detailed in Appendix C.4):
- Criterion 1: Hungary-specific: The content of the question-answer pair must align with the eight Hungary-specific dimensions proposed in this paper.
- Criterion 2: Accuracy: The information in the question-answer pair must align with the entry description and facts, and the answer should not be directly inferable from the question itself.
- Criterion 3: Concise and specific: The question and answer should be clear and concise, with no redundant information. The question should not contain nested sub-questions. The phrasing should be specific and direct, matching the scope of the answer (e.g., for time and location questions, the exact year/month/day/district/city must be specified).
- Criterion 4: Consistency Over Time: The answer should remain consistent over time and not be influenced by future events.
After the automatic checking process, we retained 5503 questions corresponding to 2666 entries.
(4) Manual review of question quality: To further ensure the quality of the questions, we hired Hungarian native speakers to manually review the questions. Annotators checked whether the questions met the four criteria mentioned in Step 3. During the annotation process, each question was assigned to two annotators, who received the questions but not the answers. A question was considered valid and retained only if both annotators agreed that it met all four criteria and that the provided answer matched the original reference answer. Detailed procedures are provided in Appendix C.4. After these four steps, we obtained a total of 1293 questions, with their distribution across the eight Hungary-specific dimensions shown in Table 16.
C.2 Obtaining corpora rich in Hungary-specific facts and knowledge
In the process of filtering Wikipedia entries with Hungarian characteristics, we randomly selected entries and provided both the entries and the first two paragraphs of the main content to GPT-4o (prompt shown in Figure 17) to determine if they were related to Hungary. If the entry was deemed relevant, it was categorized based on the eight characteristic dimensions proposed in this paper. At this stage, an “Others” category was added to ensure the focus on the eight thematic categories and to exclude interference from entries that belonged to other themes. The screening process stopped once the total number of Hungarian characteristic entries reached 8,000.
Due to the uneven distribution of entry themes on Wikipedia, with more data in the categories of figures, geography and place, and history, we filtered the data based on the proportion of themes, ensuring that no single category exceeded 1,000 entries. This resulted in 4428 characteristic entries covering the eight dimensions.
Given the varying lengths of content describing entries on Wikipedia, we aimed to streamline the complexity of constructing subsequent question-answer pairs. To achieve this, we first employed GPT-4o to extract key information from the main text of each entry. This step aims to avoid any deviation from the theme caused by redundant content during the construction of the question-answer pairs (prompt shown in Figure 18). The results of the key information extraction are presented in Figure 19.
C.3 Generating questions and corresponding answers
Based on the key information extracted and the provided entries, we utilized GPT-4o to generate 1-3 Hungarian open-ended question-answer pairs for each entry (prompt details in Figure 20). In total, 9,424 question-answer pairs were generated based on 4,000 entries.
Given that the focus and orientation of the generated question-answer pairs may differ from the original entry categories, this paper employed GPT-4o to reclassify the obtained question-answer pairs, with the corresponding prompt detailed in Figure 21.
C.4 Automatic quality checking of questions
We focused on evaluating the quality of the generated questions from two perspectives: the information contained in the question-answer pairs and the formulation of the questions. The quality assessment was divided into two stages, with each stage generating two evaluation metrics. The first stage focuses on the relevance and correctness of the question information. We provided GPT-4o with the entry, its corresponding key information, and the generated question-answer pairs to verify whether the questions contain Hungarian-specific content and whether the information in the question-answer pairs aligns with the provided background material (prompt shown in Figure 22).
Second, from the perspective of the precision of the question formulation, we only provided GPT-4o with the generated question-answer pairs to simulate real user response scenarios. This step emphasized evaluating whether the questions were based on objective facts, and whether the descriptions were precise and specific enough to allow independent answering without ambiguity. Additionally, we required that the answers remain unaffected by future events, ensuring consistency across any time period and guaranteeing the long-term validity of the dataset (prompt details in Figure 23).
Based on the results of the above automated quality assessment, we retained only those question-answer pairs that passed all four evaluation criteria, resulting in a final set of 5,503 questions.
C.5 Manual review of question quality
To further ensure the quality of the constructed question-answer pairs, we engaged native Hungarian speakers to review these questions. Each question was independently reviewed by two annotators who could only see the questions and not the reference answers. The annotation process consisted of three main steps.
First, the annotators were required to determine whether the given questions aligned with the eight Hungarian-specific dimensions proposed in this paper. Next, they evaluated whether the questions met the four assessment criteria outlined in Step 3, ensuring that the questions were objectively framed, precisely described, had unique answers, contained correct information, and maintained consistent answers over time. Finally, if a question satisfied all the above criteria, the annotators provided the correct answer. During this process, annotators were permitted to search for relevant information online and provided reference sources for their answers.
To address potential issues such as overly obscure questions or non-fixed answers, we used GPT-4o to verify whether the annotated results matched the generated reference answers. If the annotated answer matched the reference answer, it was labeled as "CORRECT"; otherwise, it was labeled as "INCORRECT" (prompt details in Figure 24). We selected question-answer pairs that both annotators deemed valid, Hungarian-specific, and consistent with the original reference answers as candidates for the HuSimpleQA dataset, resulting in a total of 2134 questions.
Considering that the HuSimpleQA dataset should exhibit diversity and broad coverage, we removed question-answer pairs belonging to the same entry, retaining only one question-answer pair per entry that best met the construction and evaluation criteria. The details of selecting optimal question-answer pairs prompt can be seen in Figure 25.
Through this process, we obtained a total of 1293 pieces of Hungarian open-ended question-answer pairs, with the category distribution shown in Table 16.
C.6 Inference prompt
We constructed prompts in two languages for model inference, as shown in Figure 26, while also instructing the model to provide a confidence score (ranging from 1 to 100)to measure the model’s confidence in its generated answers.
C.7 LLM-as-judge
Following the approach of SimpleQA Wei et al. (2024), we employed GPT-4o as the judge to evaluate the LLM’s responses. The evaluation criteria for this step were similar to those used in the manual review process of Step 4. In addition to the classification labels CORRECT and INCORRECT, we introduced an additional category, NOT_ATTEMPTED to further assess the model’s ability to respond to questions and the breadth of its knowledge coverage (prompt details in Figure 27).
Based on the results from the judge, we evaluate the performance of the LLM on HuSimpleQA using the following five metrics:
- Correct (CO): The predicted answer completely encompasses the reference answer without any conflicting or contradictory information.
- Not Attempted (NA): The predicted answer does not fully include the reference answer, but there are no contradictions between the two.
- Incorrect (IN): The predicted answer contradicts the reference answer, regardless of whether the contradiction is resolved.
- Correct Given Attempted (CGA): This metric measures the percentage of accurately answered questions out of all attempted questions.
- F-score: This metric calculates the harmonic mean between the proportion of correct answers and the proportion of correct answers among attempted questions.
The formulas for CGA and F-score are as follows:
$\text{CGA}=\frac{c}{c+i}$ (1)
[…]
Here, $c$ represents the number of CORRECTly answered questions, $i$ represents the number of INCORRECTly answered questions, and $n$ represents the number of NOT_ATTEMPTED questions.
  model correct incorrect not_attempted correct given attempted F-Score
GPT-4o 50.3 40.61 9.09 55.33 52.69
GPT-4o-mini 24.52 74.07 1.42 24.87 24.69
Deepseek-V3 32.71 64.08 3.2 33.8 33.24
QwQ 9.17 52.68 38.15 14.82 11.33
Deepseek-R1 34.58 62.15 3.28 35.75 35.15
Qwen2.5-7B-Instruct 5.29 84.13 10.58 5.92 5.59
Qwen2.5-72B-Instruct 15.05 78.61 6.33 16.07 15.54
Llama-3.1-8B-Instruct 14.9 80.25 4.84 15.66 15.27
Llama-3.1-70B-Instruct 36.36 61.03 2.61 37.34 36.84
o1-mini 16.24 44.19 39.57 26.88 20.25
Table 8: Complete results of HuSimpleQA
  Rank SimpleQA HuSimpleQA
  1 GPT-4o GPT-4o(-)
2 Deepseek-R1 Llama-3.1-Instruct-70B(↑1)
3 Llama-3.1-Instruct-70B Deepseek-R1(↓1)
4 Deepseek-V3 Deepseek-V3(-)
5 QwQ GPT-4o-mini(↑3)
6 Llama-3.1-Instruct-8B o1-mini(↑3)
7 Qwen2.5-Instruct-72B Llama-3.1-Instruct-8B(↓1)
8 GPT-4o-mini Qwen2.5-Instruct-72B(↓1)
9 o1-mini QwQ(↓4)
10 Qwen2.5-Instruct-7B Qwen2.5-Instruct-7B(-)
Table 9: Performance rankings on SimpleQA Wei et al. (2024) and HuSimpleQA
C.8 Comparing the Performance Ranking on HuSimpleQA and SimpleQA
We compare the performance rankings of LLM on HuSimpleQA and SimpleQA Wei et al. (2024), as shown in Table  9.
Appendix D HuProverbRea
D.1 Construction pipeline of HuProverb
Figure 12: Construction of HuProverbRea.
The proverbs in HuProverbRea are from 2 separate sources. The first part, 733 traditional Hungarian proverbs, are collected from the website [^14] https://mek.oszk.hu/, where each proverb is assigned an English or Hungarian explanation. The other 402 proverbs, focusing on abbreviations and Internet slang, are manually collected and explained by native speakers of Hungarian. Inspired by MAPS Liu et al. (2024c), we adopt a human-in-loop pipeline to generate and refine the context for each Hungarian specific usage, as shown in Figure 12. For each proverb, we first let GPT4 generate a seed context where the proverb is used. Then, we assign it to a Hungarian native speaker to check whether this context is grammatically correct and the use of slang is appropriate. If not, the annotator is required to manually write down a new context for the saying, which will be sent back to another annotator for inspection again. We continue the above procedures until all contexts pass the quality check. It’s worth noting that each option of the 2CQ setting is manually constructed by a human annotator, and only when it passes the double check of two other annotators could it be considered usable. We choose not to involve LLM in this part because designing correct/incorrect options requires deep understanding of sayings, LLM may generate ambiguous options if it does not understand the proverb used in the context, and such pre-provided ambiguous options may negatively influence the creativity of the annotators. Finally, we obtain 1,135 Hungarian proverbs, each equipped with a context, an English explanation, and two candidate options for question “What does the speaker mean by the saying?”.
D.2 More examples of HuProverbRea
The example of HuProverbRea is shown in Figure 28 and Figure 29. The prompt for judging HuProverbRea is shown in Figure 32.The prompt for model inference on HuProverbRea is shown in Figure 31 and 30.
  Rank HuProverbRea MAPS (en) MAPS (bn) MAPS (id)
  1 GPT-4o GPT-4o (-) GPT-4o (-) GPT-4o-mini (↑3)
2 Llama-3.1-Instruct-70B Llama-3.1-Instruct-70B (-) Deepseek-V3 (↑1) Qwen2.5-Instruct-72B (↑4)
3 Deepseek-V3 Qwen2.5-Instruct-72B (↑3) Deepseek-R1 (↑2) Llama-3.1-Instruct-70B (↓1)
4 GPT-4o-mini GPT-4o-mini (-) Qwen2.5-Instruct-72B (↑2) GPT-4o (↓3)
5 Deepseek-R1 Deepseek-V3 (↓2) o1-mini (↑2) Deepseek-V3 (↓2)
6 Qwen2.5-Instruct-72B Deepseek-R1 (↓1) GPT-4o-mini (↓2) o1-mini (↑1)
7 o1-mini Qwen2.5-Instruct-7B (↑3) Llama-3.1-Instruct-70B (↓5) Deepseek-R1 (↓2)
8 QwQ Llama-3.1-Instruct-8B (↑1) Llama-3.1-Instruct-8B (↑1) Llama-3.1-Instruct-8B (↑1)
9 Llama-3.1-Instruct-8B o1-mini (↓2) Qwen2.5-Instruct-7B (↑1) Qwen2.5-Instruct-7B (↑1)
10 Qwen2.5-Instruct-7B QwQ (↓2) QwQ (↓2) QwQ (↓2)
Table 10: Performance rankings on HuProverbRea and MAPS Liu et al. (2024c) (part1/2)
  Rank HuProverbRea MAPS (de) MAPS (ru) MAPS (zh)
  1 GPT-4o Llama-3.1-Instruct-70B (↑1) Deepseek-V3 (↑2) GPT-4o (-)
2 Llama-3.1-Instruct-70B GPT-4o-mini (↑2) o1-mini (↑5) Qwen2.5-Instruct-72B (↑4)
3 Deepseek-V3 Qwen2.5-Instruct-72B (↑3) GPT-4o (↓2) Deepseek-V3 (-)
4 GPT-4o-mini GPT-4o (↓3) Qwen2.5-Instruct-72B (↑2) o1-mini (↑3)
5 Deepseek-R1 Llama-3.1-Instruct-8B (↑4) GPT-4o-mini (↓1) Llama-3.1-Instruct-70B (↓3)
6 Qwen2.5-Instruct-72B o1-mini (↑1) Deepseek-R1 (↓1) GPT-4o-mini (↓2)
7 o1-mini Deepseek-V3 (↓4) Qwen2.5-Instruct-7B (↑3) Qwen2.5-Instruct-7B (↑3)
8 QwQ Qwen2.5-Instruct-7B (↑2) Llama-3.1-Instruct-70B (↓6) Llama-3.1-Instruct-8B (↑1)
9 Llama-3.1-Instruct-8B Deepseek-R1 (↓4) Llama-3.1-Instruct-8B (-) Deepseek-R1 (↓4)
10 Qwen2.5-Instruct-7B QwQ (↓2) QwQ (↓2) QwQ (↓2)
Table 11: Performance rankings on HuProverbRea and MAPS Liu et al. (2024c) (part2/2)
D.3 Comparing the Performance Ranking on HuProverbRea and MAPS
We compare the performance rankings of LLM on HuProverbRea and MAPS Liu et al. (2024c), as shown in Table  10 and  11.
Appendix E HuMatchingFIB and HuStandardFIB
E.1 Construction of HuMatchingFIB and HuStandardFIB
The questions for both HuMatchingFIB and HuStandardFIB are sourced from the Hungarian National Public Education Portal(NKP) [^15] https://www.nkp.hu/, a comprehensive platform for cultural funding and support in Hungary. This portal connects artists, cultural organizations, and the public with resources and opportunities to promote Hungarian culture both domestically and internationally. Notably, this website is a government initiative, reflecting the collaborative efforts between the Hungarian government and the European Union, particularly through projects or programs supported by the European Social Fund. After extracting the original questions from the NKP website, we engaged native Hungarian speakers to annotate the data. The annotation process involved manually extracting questions and their corresponding answers [^16] The questions for HuMatchingFIB and HuStandardFIB on the NKP website are not in plain text but are instead presented in interactive modules, and the answers can only be obtained through additional interactive operations. As a result, the commonly used data cleaning and extraction methods for LLM pre-training datasets are unable to accurately extract these questions and their corresponding answers. Consequently, it can be concluded that the likelihood of these questions being incorporated into the LLM pre-training data in their proper format is minimal, thereby significantly reducing the potential risk of data contamination. This ensures the reasonableness and effectiveness of the test sets for HuMatchingFIB and HuStandardFIB. , classifying the questions into appropriate categories, and filtering out questions that required additional modalities such as images, tables, audio, or video. This ensured that only purely language-based questions were retained. Through this process, we obtained 278 questions for the HuMatchingFIB task and 93 questions for the HuStandardFIB task, as shown in Table x.
E.2 More examples of HuMatchingFIB and HuStandardFIB
Examples of questions from HuMatchingFIB and HuStandardFIB are provided in Figure 33 and Figure 34. The prompt for model inference on HuMatchingFIB and HuStandardFIB is shown in Figure 36 and Figure 35.
E.3 Metric and Judge of HuMatchingFIB and HuStandardFIB
HuMatchingFIB employs a rule-based evaluation approach, where the assessment is conducted at two levels: the blank level and the question level (as a single question may contain multiple blanks). The evaluation process is analogous to that of multiple-choice questions, and accuracy (acc) is used as the metric to determine performance. The corresponding formula for blank level accuracy is as follows, where c represents the number of correctly predicted blanks in one question, t represents the number of blanks in one question.
$\text{Acc}_{\text{blank level}}=\frac{\sum\text{blank}_{\text{c}}}{\sum\text{blank}_{\text{t}}}$ (3)
HuStandardFIB questions are designed with open-ended reference answers to accommodate variations in part of speech and semantics. We employ a many-to-one fuzzy matching mechanism. Fuzzy matching is a technique that calculates the similarity between strings, allowing for flexibility in matching by considering variations such as typos, synonyms, or different word orders. In this context, the model’s answer is compared against a set of possible reference answers (where multiple correct answers may exist for a single question or blank). If the similarity score between the model’s answer and any of the reference answers exceeds a predefined threshold, the answer is considered correct. This approach is particularly suitable for evaluating open-ended questions where exact matches are often infeasible due to the variability in acceptable responses. The annotator information involved in all tasks of OpenHuEval can be found in Appendix G.
Appendix F Analyzing LRM’s thinking process on OpenHuEval
F.1 Analyzing LRM’s thinking process on HuSimpleQA
We use GPT-4o to break down the answers into thoughts. This is done in two steps: the first step is to identify expressions that may be a shift in thought (see the prompt in Figure 37), and the second step is to confirm whether it is indeed a shift in thought (set the prompt in Figure 38). Then, We utilized the LLM to evaluate whether each idea would lead to the correct answer, the prompt is shown in Figure 39. We consider a confident score of 2 as the correct thought.
Examples of the thoughts and corresponding correctness can be found in Figure 40 and Figure 41.
F.2 Analyzing LRM’s thinking process on HuMatchingFIB
F.2.1 Splitting LRM’s thinking process into segments and classifying these segments
We first divided the thinking process of LRMs into multiple segments, with each segment categorized as introduction, reasoning, review, or summary. The definitions of these four segment types are provided in Table 12. Typically, each thinking process begins with an introduction segment, includes several reasoning segments and some review segments in the middle, and concludes with a summary segment. The segmentation and classification were performed using GPT-4o, with the prompt template detailed in Figures 42, 43 and 44. Examples of the segmentation and classification can be found in Figure 47 and Figure 48.
  Segment types Definition
introduction typically located at the beginning of the reasoning process; usually consists of the LRM’s brief restatement of the question and the descriptive account of the work it is about to undertake; does not include the actual start of the analysis of the question.
reasoning typically constitutes the main body of the reasoning process; includes the detailed thinking and reasoning steps undertaken by the LRM to solve the fill-in-the-blank questions.
review usually occurs after the reasoning process is essentially complete but before the final output. This section typically includes a review of the entire reasoning process and may contain keywords or phrases such as “Overall, …” or “double check…” .
summary Summarizes the overall content or provides final conclusions, often using phrases like ’in conclusion’ or ’overall’.
Table 12: Definitions of the four categories of segments in the thinking process of LRM on HuMatchingFIB.
F.2.2 Tagging the reasoning segments along the dimensions
We further tagged the reasoning segments according to the following four dimensions: (Dim1) correctness: Are the answers in this reasoning segment correct? (Dim2) complexity: In this reasoning segment, does the LRM simply assert the answer, or does it involve more complex reasoning? (Dim3) scope: Does this reasoning segment focuses on a single blank, modifies previous blanks, or addresses multiple blanks? (Dim4) language transfer: Does the LRM switch languages within this reasoning segment? The details of the tagging can be found in Table 13. Examples of the tagging results can be found in Figure 47 and Figure 48.
  Dimension Description Tags
Dim 1: Correctness Are the answers in this reasoning segment correct? Class 4 is used when no conclusion is reached.
Class 1: Completely Incorrect
Class 2: Partially Correct
Class 3: Completely Correct
Class 4: Non Conclusion
Dim 2: Complexity In this reasoning segment, does the LRM simply assert the answer, or does it involve more complex reasoning?
Class 1: Simple Assertion
Class 2: Complex Thought
Dim 3: Scope Does this reasoning segment focuses on a single blank, modifies previous blanks, or addresses multiple blanks?
Class 1: Only Current Blank
Class 2: Modify Previous Blanks
Class 3: Current Blank and Consecutive Blank
Dim 4: Language Transfer Does the LRM switch languages within this reasoning segment (e.g., Hungarian to English)?
Class 1: Contains Language Transfer
Class 2: No Language Transfer
Table 13: Tagging dimensions of the reasoning segments in LRM’s thinking process on HuMatchingFIB
Appendix G Information of the Annotators
  Task # Anotater Total working hours
HuSimpleQA 14 161.9
HuWildBench 5 55.2
HuProverbRea 15 118.2
HuMatchingFIB and
HuStandardFIB
8 84.5
Table 14: Information of the Annotators
We submitted the annotation task online to a professional data annotation company, which organized annotators to complete the annotation work. In the construction phase of OpenHuEval, the annotations were carried out by professional annotators who are native Hungarian speakers. Table 14 shows the number of annotators and the total time spent on each task. All annotators involved in this project hold a bachelor’s or master’s degree, with academic backgrounds in fields such as Social Sciences, Translating and Interpreting, English Studies, and IT Engineering. They all possess the ability to distinguish subtle aspects of the Hungarian language and handle Hungarian-specific knowledge effectively.
Figure 13: Prompt template for automatic filtering of user questions related to Hungarian specifics in the construction of HuWildBench.
Figure 14: Prompt template for automatic filtering of high-quality question in the construction of HuWildBench.
  Hungarian-specific dimensions: $\mathcal{LCC}$ Count: 365
Question Example:
a kérdés az: Mi lesz a jövőben a szocializmus alatt megépül sok panellel?
a leírás: Úgy tudom, hogy kb 60 éves életciklusra tervezték őket. Magyarországon (és a környező országokban is) rengeteg ember él bennük. Mi fog történni akkor, ha lakhatatlanná kezdenek válni? Mi lesz azzal a sok emberrel? Meg a panelokkal?
Translation to EN:
The question is: What will happen in the future to the many panels built under socialism?
The description is: I understand they are designed for a life cycle of about 60 years. There are a lot of people living in them in Hungary (and surrounding countries). What will happen if they start to become uninhabitable? What will happen to all those people? And the panels?
Checklist:
"Does the response provide an analysis of the current condition and expected lifespan of the panel buildings in Hungary and neighboring countries?",
"Does the response address the expected lifespan of panel buildings and their current age?",
"Are there any historical or international examples included to illustrate possible outcomes or strategies?",
"Does the response consider the economic implications of renovating or replacing panel buildings?",
"Does the response include potential government or private sector plans or policies addressing the future of these buildings and their residents?",
"Does the answer discuss potential scenarios for when these buildings become uninhabitable?",
" Are environmental and urban planning aspects of dealing with aging panel buildings mentioned?",
"Is there an explanation of possible solutions or government plans for relocating residents?"
  Hungarian-specific dimensions: $\mathcal{EP}$ Count: 201
Question Example:
a kérdés az: A kárpátaljai magyarok Ukrajnában oroszul vagy ukránul tanultak meg a 2000-es évek közepén?
a leírás: Mit tanítottak az iskolákban? Mennyire reális az, hogy valakire szinte semmi se ragad a környezetéből? Vannak olyan tömb területek ahol mondjuk egy magyar gyereknek egyáltalán nem kell helyi ukránokkal beszélnie? Egyáltalán a helyi ukránok ukránul beszéltek a 2000-es években?
Translation to EN:
The question is: Did Hungarians in Transcarpathia learn Russian or Ukrainian in Ukraine in the mid-2000s?
The description is: What was taught in schools? How realistic is it that almost nothing sticks to someone from their environment? Are there block areas where, say, a Hungarian child doesn’t have to speak to local Ukrainians at all? Did local Ukrainians even speak Ukrainian in the 2000s?
Checklist:
"Does the answer provide information on the language predominantly spoken by local Ukrainians in Transcarpathia in the 2000s?",
"Does the response discuss the social and linguistic dynamics in areas with significant Hungarian populations, including interactions with local Ukrainians?",
"Does the response clearly explain the educational policies and language of instruction in schools for Hungarians in Transcarpathia during the mid-2000s?",
"Does the response accurately describe the language of instruction in Transcarpathian Hungarian schools in the mid-2000s?",
"Does the response consider the historical and political context of language policies in Ukraine during this period?",
"Does the response provide insight into whether local Ukrainians predominantly spoke Ukrainian during the 2000s?",
"Does the response offer a balanced view of cultural and linguistic integration in Transcarpathia during the specified period?",
"Does the answer address the likelihood of a Hungarian child not acquiring any local language skills from their environment?",
"Does the response discuss the existence of predominantly Hungarian areas where interaction with local Ukrainians might be limited?"
  Hungarian-specific dimensions: $\mathcal{PPL}$ Count: 299
Question Example:
a kérdés az: Mi történt azzal, aki az 50-es években a felhívás ellenére sem jegyzett "önként" békekölcsönt? Érhette ezért retorzió az embert?
a leírás: Persze nyilván volt, amilyen "bolondos" idők jártak nálunk akkortájt. Biztos kikiáltották reakciósnak vagy fasisztának, meg a "népi demokrácia" ellenségének.
Translation to EN:
The question is: What happened to the man who did not "voluntarily" subscribe to a peace charter in the 1950s, despite the call? Could he have been retaliated against for this?
The description is: He must have been branded a reactionary or a fascist or an enemy of ’people’s democracy’.
Checklist:
"Does the answer address the political labels mentioned in the description (e.g., ’reactionary’, ’fascist’, ’enemy of people’s democracy’)?",
"Does the response differentiate between official consequences and social/societal repercussions for not subscribing to the peace loan?",
"Does the response address potential consequences for individuals who did not subscribe to the peace loan, with references to historical examples or documentation?",
"Does the response provide a balanced view, considering both potential punitive measures and any instances of leniency or exceptions, if applicable?",
"Is there a clear explanation of what ’békekölcsön’ (peace loan) was and its significance during that time period?",
"Does the response accurately describe the historical context of the 1950s in Hungary?",
"Is there an analysis of the societal and governmental attitudes toward dissenters in Hungary during the 1950s, including any possible labels or accusations they might have faced?",
"Does the response provide specific examples of potential retaliations against those who didn’t subscribe to the peace loan?"
  Hungarian-specific dimensions: $\mathcal{BF}$ Count: 289
Question Example:
a kérdés az: Meddig tartható fent Magyarország negatív külkereskedelmi mérlege?
a leírás: Nem a háború óta, hanem már 2021 nyarától folyamatosan negatív az ország külkereskedelmi mérlege. Júliusban és augusztusban összesen több, mint 1000 milliárd forintnyi mínusz keletkezett. Persze a többi hónap nem volt ennyire szörnyű, de ez csak erre az évre már több, mint 2000 milliárd forintnyi mínusz. Változatlan devizaimport mellett a mérséklődött energiaárakkal is több, mint 1000 milliárdos negatív mérleg hozható össze 2023-ban. Meddig lehet ezt tovább folytatni? Meddig elég a devizatartalék a hiány pótlására?
Translation to EN:
The question is: How long can Hungary maintain a negative trade balance?
The description is: In July and August there was a total deficit of more than HUF 1000 billion. Of course, the other months were not so bad, but for this year alone it is already more than HUF 2000 billion in deficit. Even with unchanged foreign exchange imports and moderating energy prices, a negative balance of more than 1,000 billion in 2023 could be created. How long can this go on? How long will foreign exchange reserves be enough to cover the deficit?
Checklist:
"Does the response analyze Hungary’s current foreign exchange reserves and their sufficiency in covering the trade deficit?",
"Is there an exploration of historical trends and comparisons to similar situations in other countries to provide context?",
"Is the impact of energy prices on the trade balance accurately assessed in the response?",
"Does the response offer a clear and supported prediction or timeframe for how long Hungary can sustain its negative trade balance?",
"Is there an analysis of the factors affecting Hungary’s foreign exchange reserves and their ability to cover the deficit?",
"Does the answer provide a clear timeline or projection for how long the negative balance can be sustained?",
"Are there comparisons made to similar situations in other countries or historical precedents in Hungary?",
"Does the response accurately explain the current state of Hungary’s foreign trade balance?"
Table 15: Examples of HuWildBench. The blue font is the English translation of the original OpenHuEval examples, used for visualization.
Figure 15: Prompt template for constructing the checklist in the construction of HuWildBench.
Figure 16: Prompt template for LLM as judge on HuWildBench (WBScore).
Figure 17: Prompt template for selecting Hungarian specific Wikipedia entries in the construction of HuSimpleQA.
Figure 18: Prompt template for extracting key information from the wikipedia entries in the construction of HuSimpleQA.
Figure 19: Example of the extracted key information from the Hungarian wikipedia entries. (HuSimpleQA)
Figure 20: Prompt template for generating Hungarian question-answer pairs in the construction of HuSimpleQA.
Figure 21: Prompt template for categorizing the generated question-answer pairs in the construction of HuSimpleQA.
Figure 22: Prompt template for evaluating the relevance and correctness of question-answer pairs in the construction of HuSimpleQA.
Figure 23: Prompt template for evaluating the precision and consistency of question-answer pairs in the construction of HuSimpleQA.
Figure 24: Prompt template for evaluating human-annotated answers in the construction of HuSimpleQA.
Figure 25: Prompt template for selecting optimal question-answer pairs in the construction of HuSimpleQA.
Figure 26: Prompt template for model inference on HuSimpleQA.The left is the original prompt, the right is the English translation for visualization
Hungarian-specific
dimensions
Count Question-Answer Pairs translation to EN
  L 11
Question1: Mit jelent a Kara török eredetű régi magyar személynév?
Answer1: fekete
Question2: Melyik régi magyar név a Pantaleon megfelelője?
Answer2: Pentele
Question1: What does the old Hungarian personal name Kara of Turkish origin mean?
Answer1: black
Question2: Which old Hungarian name is the equivalent of Pantaleon?
Answer2: Pentele
  H 140
Question1: Melyik király nevezte ki Szapolyai Imrét szepesi örökletes főispánná 1465-ben?
Answer1: Mátyás király
Question2: Melyik várost foglalta el Báthory Gábor 1610. december 11-én?
Answer2: Szeben
Question1: Which king appointed Imre Szapolyai as the hereditary ispán of Szepes in 1465?
Answer1: King Matthias
Question2: Which city was captured by Gabriel Báthory on December 11, 1610?
Answer2: Sibiu
  LCC 257
Question1: Melyik magyar film nyerte el a FIPRESCI-díjat az 1983-as Cannes-i Nemzetközi Filmfesztiválon?
Answer1: Szerencsés Dániel
Question2: Melyik legendára épít az ’Eredet / Origins’ táncjáték?
Answer2: Csodaszarvas-legendára
Question1: Which Hungarian film won the FIPRESCI Prize at the 1983 Cannes International Film Festival?
Answer1: Lucky Daniel
Question2: Which legend is the ’Origin / Origins’ dance play based on?
Answer2: Legend of the Miraculous Deer
  EP 81
Question1: Melyik városban alapították a Gandhi Gimnáziumot 1994-ben?
Answer1: Pécsen
Question2: Melyik évben alapította a Magyar Tudományos Akadémia az Acta Juridica Hungarica folyóiratot?
Answer2: 1959
Question1: In which city was the Gandhi High School founded in 1994?
Answer1: Pécs
Question2: In which year did the Hungarian Academy of Sciences establish the journal Acta Juridica Hungarica?
Answer2: 1959
  GP 165
Question1: Melyik magyar vármegyében található Nemesmedves?
Answer1: Vas vármegyében
Question2: Mi a neve Magyarország legmagasabban fekvő csillagvizsgálójának, amely a Piszkés-tetőn található?
Answer2: Piszkéstetői Obszervatórium
Question1: In which Hungarian county is Nemesmedves located?
Answer1: Vas county
Question2: What is the name of Hungary’s highest observatory, located on Piszkés Peak?
Answer2: Piszkés Peak Observatory
  F 409
Question1: Nádasdy Kálmán hányszor kapott Kossuth-díjat élete során?
Answer1: Háromszor
Question2: Balogh József melyik magyar városban született 1946. április 15-én?
Answer2: Nagykanizsán
Question1: How many times did Kálmán Nádasdy receive the Kossuth Prize during his lifetime?
Answer1: Three times
Question2: In which Hungarian city was József Balogh born on April 15, 1946?
Answer2: Nagykanizsa
  PPL 186
Question1: Melyik szervezet jogkörét vette át a Népgazdasági Tanács 1949. június 11-én?
Answer1: Gazdasági Főtanács
Question2: Melyik törvénycikk rendelkezett 1878-ban Magyarországon a réz-váltópénz szaporításáról?
Answer2: 1878. évi VI. törvénycikk
Question1: Which organization’s authority was taken over by the National Economic Council on June 11, 1949?
Answer1: Supreme Economic Council
Question2: Which statute regulated the increase of copper coinage in Hungary in 1878?
Answer2: Act VI of 1878
  BF 44
Question1: Milyen néven működött az ÉVITERV 1954-től az 1980-as évek elejéig?
Answer1: ÉM Szerelőipari Tervező Vállalat
Question2: Melyik cég gyártotta a Puli autótípust a gyártás kezdeti időszakában?
Answer2: HÓDGÉP
Question1: Under what name did ÉVITERV operate from 1954 to the early 1980s?
Answer1: ÉM Installation Industry Design Company
Question2: Which company manufactured the Puli car model in the early production period?
Answer2: HÓDGÉP
Table 16: Examples of HuSimpleQA. The rightmost column is the English translation of the original OpenHuEval examples, used for visualization.
Figure 27: Prompt template for LLM as judge on HuSimpleQA.
Figure 28: Example of HuProverbRea (2CQ). The left is the original example in OpenHuEval, the right is the English translation for visualization.
Figure 29: Example of HuProverbRea (OE). The left is the original example in OpenHuEval, the right is the English translation for visualization.
Figure 30: Prompt template for model inference on HuProverbRea (OE).
Figure 31: Prompt template for model inference on HuProverbRea (2CQ).
Figure 32: Prompt template for LLM as judge on HuProverbRea.
Figure 33: Example of HuMatchingFIB. The left is the original example in OpenHuEval, the right is the English translation for visualization.
Figure 34: Example of HuStandardFIB. The left is the original example in OpenHuEval, the right is the English translation for visualization.
Figure 35: Prompt template for model inference on HuStandardFIB.
Figure 36: Prompt template for model inference on HuMatchingFIB.
Figure 37: Prompt template for splitting the LRM’s thinking process into thoughts on HuSimpleQA (step 1/2).
Figure 38: Prompt template for splitting the LRM’s thinking process into thoughts on HuSimpleQA (step 2/2).
Figure 39: Prompt template for evaluating the correctness of the thoughts of the LRM’s thinking process on HuSimpleQA.
Figure 40: Example of the thoughts in Deepseek-R1’s thinking process on HuSimpleQA. and $\times$ indicate the correctness of the thoughts. The original text is in black, while the translation into English is in blue.
Figure 41: Example of the thoughts in QWQ’s thinking process on HuSimpleQA. and $\times$ indicate the correctness of the thoughts. The original text is in black, while the translation into English is in blue.
Figure 42: Prompt template for splitting LRM’s thinking process into segments and categorizing these segments on HuMatchingFIB. (part 1/3).
Figure 43: Prompt template for splitting LRM’s thinking process into segments and categorizing these segments on HuMatchingFIB. (part 2/3).
Figure 44: Prompt template for splitting LRM’s thinking process into segments and categorizing these segments on HuMatchingFIB. (part 3/3).
Figure 45: Prompt template for tagging the reasoning segments of LRM’s thinking process along the four dimensions on HuMatchingFIB (part 1/2).
Figure 46: Prompt template for tagging the reasoning segments of LRM’s thinking process along the four dimensions on HuMatchingFIB (part 2/2).
Figure 47: Example of splitting Deepseek-R1’s thinking process into segments, categorizing these segments, and tagging the reasoning segments on HuMatchingFIB. The categorizing results are in green and the tagging results are in red.
Figure 48: Example of splitting QwQ’s thinking process into segments, categorizing these segments, and tagging the reasoning segments on HuMatchingFIB. The categorizing results are in green and the tagging results are in red.